OpenStudy (anonymous):

how we can calculate magnetic field .when The electric pass through The wire .

OpenStudy (anonymous):

/\[\mu_0*i*(dl \times R)/4*\pi*r^3\] where, r=mod(R)

OpenStudy (ybarrap):

Use Ampere's law, without Maxwell's displacement term (since this is zero anyway)